OpenStudy (anonymous):

Select the pair of equations whose graphs are perpendicular. 2x – 8y = 9 12x – 3y = 7 3x + 6y = 8 y = 2x – 8 y = 2x – 7 x + y = 3 2y = –3x + 5 2x + 3y = 4

Directrix (directrix):

The slopes of perpendicular lines have a product of negative one.

OpenStudy (anonymous):

Firstly put eq in y = mx+c form m = slope for a line to be perpendicular m(perpendicular) = 1/(-m) of the given line

Directrix (directrix):

3x + 6y = 8 --> 6y = 8 -3x, slope is -3/6 = -1/2* y = 2x – 8 --> slope is 2/1 Question: What is the product of (-1/2) * (2/1) ? @CrazyCountryGirl

OpenStudy (anonymous):

@Directrix -1

Directrix (directrix):

Yes, so using the theorem that the product of the slopes of perpendicular lines is -1, provided neither line is horizontal or vertical, you can deduce that these two lines of the second option are perpendicular. As shown, the slopes of these two lines have a product of -1. @CrazyCountryGirl 3x + 6y = 8 y = 2x – 8